Subject: APO 3.5m users committee minutes, 1/31/11
From: Suzanne Hawley
Submitted: Fri, 11 Feb 2011 11:04:57 -0800 (PST)
Message number: 1216
(previous: 1215,
next: 1217
up: Index)
APO 3.5-m Users Committee Phonecon, 1/31/11
Attending: Suzanne Hawley, Mark Klaene, Michael Strauss, Bill Ketzeback,
Russet McMillan, Remy Indebetouw, Sean Moran, Jon Holtzman, Rene Walterbos,
Bruce Gillespie
Absent: Scott Anderson, John Bally, Al Harper
**********************************
User feedback, comments from institutional representatives:
o Chicago (Al Harper) - no report.
o New Mexico State (Jon Holtzman) - Jon had nothing to report.
o Washington (Suzanne Hawley) - Suzanne (for Scott) had nothing to report.
o Colorado (John Bally) - no report.
o Princeton (Michael Strauss) - Michael said the PU users are happy.
o Johns Hopkins (Sean Moran) - Sean had nothing to report.
o Virginia (Remy Indebetouw) - Remy said that his users are generally
happy. There has been an issue with synchronizing the guider images and
spectra from Triplespec. The names in the file headers seem to be off by
one image. Remy agreed to ask Matt Nelson at UVa about how the headers
are written and report back.
**********************************
Discussion of telescope/instruments report - Klaene:
Mark mentioned the forecast for the first major winter storm this winter,
expected to hit us this week. The matplotlib memory leak mentioned below
has been mitigated, but won't be eliminated until the software vendor
fixes their code. The telescope has generally behaved well, but pointing
errors are slowly creeping upward. On the instruments, the condensation
on the SPIcam chip has been removed by heating the detector and pumping --
we are running the chip heater whenever the instrument is not being used
for observing. Agile is experiencing spontaneous aborts which is due to
the software drivers from Princeton Instruments. TripleSpec has been
occasionally giving zero-filled frames, and resetting the controller seems
to fix it temporarily. Low-level banding seen in echelle images was
looked into by Dale Sandford from Yerkes, one of the developers. No
solution is yet obvious, but it seems the problem may be related to
instrument and/or telescope grounding. The Goddard Fabry-Perot made its
final run prior to its upgrade that begins next week. The echelle shutter
is not functional, and is being investigated today. Mark's detailed
report for the reporting period follows:
*************************
3.5-m Telescope, Instruments, and CIF Projects Highlights
12/2/10 through 1/26/11
0) Overview
Weather turned slightly worse this period with winter storms bringing
clouds but very little precipitation. Cloud cover kept us closed a few
whole nights but many nights we were able to open for a few hours. We
also managed to get through the holidays with very few issues that
required day staff to come in during their break.
We started filming for the new APO promotional video and "B-roll" footage
in high definition. The primary effort was during the recent lunar eclipse
when we acquired a couple very cool images of the APOLLO laser in
operation.
1) Telescope
Virtually no telescope problems were reported during this period. We are
seeing an occasional aborting of some of the guiders (Na2 and TCam) that
is being investigated. The restart is quick and the new TUI guiding plots
make noticing the aborts much easier. New TUI 1.8.5 and 1.8.6b versions
were issued. Note - memory leak is still present but at a much reduced
level. Changes are:
- Updated for backward-incompatible StripChartWdg
- The Mac version now includes matplotlib 1.0.1 and unix users may
wish to upgrade as well. This offers two important benefits: TUI will no
longer exit at startup if matplotlib's font cache is outdated; and
Matplotlib's memory leak in canvas.draw is reduced by a factor of 10.
- Added target axis position and UTC date to the Status window.
- Added Calibration offsets to the Nudger window.
- Renamed Object offsets to Object Arc offsets in the Offset,
Nudger and Status windows and reorganized the associated help files to
better consolidate information about types of offsets.
- Enclosure window: added eyelids summary status (which tells you
if the eyelid for the current port is open) and Open All and Close All
buttons.
- fixed issue with unwanted ctrl-click in slitviewer windows.
2) Instruments
DIS has performed well this period with only minor issues with the
slit-viewer shutter after its recent replacement.
SPIcam was discovered to have some condensation on the detector again.
After last month's vacuum pumping we noticed that the CCD was running 10
deg colder than normal when the condensation appeared. We ran the chip
heater to get to the nominal setpoint and the condensation then
disappeared. This allowed us to continue observations for the time being.
The chip heater when running at greater than 50% does induce some
artifacts so the heater is shutoff prior to observing. We are
investigating a current limit such that the heater could stay on
continuously at a low level. Agile spontaneous aborts are being
investigated. Despite repeated attempts, the camera vendor (Princeton
Instruments) does not seem to be able to offer a solution for this in
Linux. We are investigating some software updates and a possible move to
a Windows platform. A decision will be made after this week's observing
as to the seriousness of the problem.
TripleSpec is not reporting slit position in the headers. This is on the
list of things to fix.
Echelle continued showing some banding in bias images. Dale Sandford from
Yerkes was on site (week of Jan 24-28) troubleshooting this. The problem
does show up in the lab at about a 1/2 count level but not repeatedly,
much like at the telescope.
NIC-FPS was returned to CU to address the readout problem when rotated.
We also hope the new fanout board will be installed during this trip.
GFP had a design review for the new detector and the IFS and will be taken
off line after their run next week to start the upgrade. Initial
commissioning using their control system is scheduled for March.
**********************************
NIC-FPS update - Klaene:
The instrument is currently at Colorado, Mark said. The good news is that
the new fan-out board was installed and has cut the read noise by a factor
of two, almost to the level we saw before the 16-channel read conversion
was done several years ago. We are still expecting the instrument to
return to APO in time for the March bright run. The problem with the
instrument malfunctioning seen when rotating has yet to be diagnosed.
**********************************
TUI for windows machines - Hawley:
Suzanne said that Mac OS10 v.10.4 will soon be incompatible with the
current version of TUI. There is also the question for TUI users who use
Windows machines -- maintaining TUI for the older version of Mac OS and
Windows is becoming very difficult, so we need to assess our need for
supporting TUI for Windows platforms:
===> ACTION: Users Committee members should poll their users to find out
how many are using TUI with Windows machines and report at the next Users
Committee meeting.
**********************************
ReSTAR proposal - Holtzman, Hawley:
ReSTAR Phase 2 is starting, and Jon attended a kick-off meeting in Tucson
in November. The pre-proposal paper that we crafted in the past weeks
will be submitted tomorrow. This is to be construed as a "foot in the
door", and enables NOAO to work with us to negotiate observing time for
instrument(s) and/or cash, and other forms of short and long-term support
and collaboration, if we choose to go this route. One thing we need to
address more fully is our prioritization of what we want our
instrumentation suite to evolve into over the coming years.
**********************************
Class visits - Klaene:
For everyone's information, there is a worsening housing shortage at
NSO-Sunspot due to the many staff they are hiring for their ATST project.
This means for example that some SDSS observers have been displaced and
now have to be in the dorms while they are observing. In mid-March, we
are expecting a class from Colorado with up to 19 students and faculty;
this will be hard to accommodate between the available housing at APO and
Sunspot, and other liens in the period for APO housing. For future
reference, we believe that class visits of less than 10 people total would
be the preferable group size, at least for the next few years.
**********************************
ACTION ITEMS:
[open from previous meeting] -- none
[new actions from this meeting]:
===> ACTION: Users Committee members should poll their users to find out
how many are using TUI with Windows machines and report at the next Users
Committee meeting.
**********************************
Next meeting: The next Users Committee phonecon will be on Monday 28
February at 8:30 AM Pacific Time. The agenda and other materials will be
sent to the committee members during the preceding week.
APO APO APO APO APO Apache Point Observatory 3.5m APO APO APO
APO
APO This is message 1216 in the apo35-general archive. You can find
APO the archive on http://www.astro.princeton.edu/APO/apo35-general/INDEX.html
APO To join/leave the list, send mail to apo35-request@astro.princeton.edu
APO To post a message, mail it to apo35-general@astro.princeton.edu
APO
APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO